OpenStudy (anonymous):

hi my name is alicia and am havin truble finding the gcf of 252 and 480

OpenStudy (nikvist):

\[252=2\cdot 2\cdot 3\cdot\ 3\cdot 7\] \[480=2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 3\cdot 5\] GCF is \[2\cdot 2\cdot 3=12\]

OpenStudy (anonymous):

480 - 252 = 228 252 - 228 = 24 252 - 10 x 24 = 12 12 | 24 => 12 is the GCF See 'Euclidean Algorithm' for more. You can do it the babyish way above for small numbers, but when prime factorisation is computationally difficult, you'll come crawling back to Euclid.